# On the Graphical Rules for Recovering the Average Treatment Effect Under Selection Bias

### v4

@@ 1 Introduction
With a clearer understanding of selection bias in place, recent methodological developments have been dedicated to recovering causal effects in the presence of selection bias. Specifically, aided by causal diagrams, including directed acyclic graphs (DAGs) and single-world intervention graphs (SWIGs)[richardson2013single, bezuidenhout2025swigs], several graphical rules were proposed to assess and account for selection bias, either alone or in combination with other challenges (e.g., confounding bias or missing data) [bareinboim2014recovering, correa2018generalized, mathur2023simple, schnell2024graphical].
While these graphical rules provide valuable insights, they are typically paired with specific identification strategies or estimators. For example, under the graphical rules proposed by Mathur and Shpitser, one key premise is that post-treatment variables cannot be conditioned on when conducting selected-sample analysis (i.e., complete-case analysis) [mathur2023simple]. In addition, their focus is on the conditional average treatment effect (CATE), rather than the average treatment effect (ATE) which is arguably more widely used. By contrast, epidemiologic analyses have frequently relied on inverse probability of selection or censoring weights to address different forms of selection bias (e.g., due to non-random sample selection or differential loss to follow-up). [hernan2004structural, howe2016selection] These approaches leverage covariate information (potentially including post-treatment variables) from unselected or censored individuals, rather than relying solely on selected-sample analysis. Consequently, the graphical rules proposed by Mathur and Shpitser may not be universally applicable. Other existing graphical rules, while not prohibiting adjustment for post-treatment variables, impose stringent restrictions intended to avoid confounding bias [correa2017causal, correa2018generalized, bareinboim2014recovering].
@@ 2 Two cases of selection bias unaddressed by simple graphical rules
To detect and control selection bias, researchers often rely on simple graphical rules in conjunction with a specific class of identification formulas, and define identifiability (or recoverability) narrowly as whether the treatment effect of interest can be identified within that class [correa2017causal, correa2018generalized, bareinboim2014recovering, mathur2023simple]. These graphical rules typically provide sufficient conditions for establishing that the treatment effect is identifiable (or recoverable) using a given set of formulas. However, failure to satisfy these conditions does not imply that the effect is unidentifiable; rather, it indicates that identification may require alternative strategies or estimators. In this section, we present two important cases of selection bias that fall outside the scope of previously proposed graphical rules. Identification of treatment effects for the related causal structures using modified formulas is discussed in later sections.
@@ 2.2 Limitations of existing simple graphical rules
We focus on three existing graphical rules proposed over the past decade: the selection-backdoor criterion [bareinboim2014recovering], the generalized adjustment criterion [correa2018generalized], and the graphical rules of Mathur and Shpitser [mathur2023simple]. In Appendix A, we demonstrate that the two cases discussed above violate these rules, and that identification by their corresponding formulas is therefore not guaranteed.
@@ 6 Discussion
Our work assesses and extends the utility of existing graphical rules for addressing selection bias. Derived based on the fundamental principle of reading conditional independence using d-separation, existing simple graphical rules are inherently coupled with specific identification formulas and are therefore applicable only to a subset of selection mechanisms. We show that, some important existing rules [bareinboim2014recovering, correa2018generalized, mathur2023simple] can impose stringent restrictions on covariate adjustment in order to avoid confounding bias while correcting for selection bias. For example, adjustment for post-treatment variables is generally discouraged under these rules, even in settings where such variables help separate selection from the outcome. While post-treatment variables should not be adjusted for when addressing confounding bias, they can nonetheless be appropriately accounted for when dealing with selection bias using methods such as IPW or the g-formula. Using two concrete examples illustrated with causal diagrams, we demonstrate the limitations of existing graphical rules and their associated identification strategies for addressing selection bias.
@@ A.3 Simple graphical rules under SWIG by Mathur and Shpitser
Graphical rules based on SWIGs were also proposed for addressing selection bias [mathur2023simple]. It involves two sufficient conditions to analyze whether a conditional average treatment effect (CATE) can be identified by their formula.
Definition 3 (The graphical rules for assessing selection bias [mathur2023simple]).

### v5

@@ 1 Introduction
With a clearer understanding of selection bias in place, recent methodological developments have been dedicated to recovering causal effects in the presence of selection bias. Specifically, aided by causal diagrams, including directed acyclic graphs (DAGs) and single-world intervention graphs (SWIGs)[richardson2013single, bezuidenhout2025swigs], several graphical rules were proposed to assess and account for selection bias, either alone or in combination with other challenges (e.g., confounding bias or missing data) [bareinboim2014recovering, correa2018generalized, mathur2025simple, schnell2024graphical].
While these graphical rules provide valuable insights, they are typically paired with specific identification strategies or estimators. For example, under the graphical rules proposed by Mathur and Shpitser, one key premise is that post-treatment variables cannot be conditioned on when conducting selected-sample analysis (i.e., complete-case analysis) [mathur2025simple]. In addition, their focus is on the conditional average treatment effect (CATE), rather than the average treatment effect (ATE) which is arguably more widely used. By contrast, epidemiologic analyses have frequently relied on inverse probability of selection or censoring weights to address different forms of selection bias (e.g., due to non-random sample selection or differential loss to follow-up). [hernan2004structural, howe2016selection] These approaches leverage covariate information (potentially including post-treatment variables) from unselected or censored individuals, rather than relying solely on selected-sample analysis. Consequently, the graphical rules proposed by Mathur and Shpitser may not be universally applicable. Other existing graphical rules, while not prohibiting adjustment for post-treatment variables, impose stringent restrictions intended to avoid confounding bias [correa2017causal, correa2018generalized, bareinboim2014recovering].
@@ 2 Two cases of selection bias unaddressed by simple graphical rules
To detect and control selection bias, researchers often rely on simple graphical rules in conjunction with a specific class of identification formulas, and define identifiability (or recoverability) narrowly as whether the treatment effect of interest can be identified within that class [correa2017causal, correa2018generalized, bareinboim2014recovering, mathur2025simple]. These graphical rules typically provide sufficient conditions for establishing that the treatment effect is identifiable (or recoverable) using a given set of formulas. However, failure to satisfy these conditions does not imply that the effect is unidentifiable; rather, it indicates that identification may require alternative strategies or estimators. In this section, we present two important cases of selection bias that fall outside the scope of previously proposed graphical rules. Identification of treatment effects for the related causal structures using modified formulas is discussed in later sections.
@@ 2.2 Limitations of existing simple graphical rules
We focus on three existing graphical rules proposed over the past decade: the selection-backdoor criterion [bareinboim2014recovering], the generalized adjustment criterion [correa2018generalized], and the graphical rules of Mathur and Shpitser [mathur2025simple]. In Appendix A, we demonstrate that the two cases discussed above violate these rules, and that identification by their corresponding formulas is therefore not guaranteed.
@@ 6 Discussion
Our work assesses and extends the utility of existing graphical rules for addressing selection bias. Derived based on the fundamental principle of reading conditional independence using d-separation, existing simple graphical rules are inherently coupled with specific identification formulas and are therefore applicable only to a subset of selection mechanisms. We show that, some important existing rules [bareinboim2014recovering, correa2018generalized, mathur2025simple] can impose stringent restrictions on covariate adjustment in order to avoid confounding bias while correcting for selection bias. For example, adjustment for post-treatment variables is generally discouraged under these rules, even in settings where such variables help separate selection from the outcome. While post-treatment variables should not be adjusted for when addressing confounding bias, they can nonetheless be appropriately accounted for when dealing with selection bias using methods such as IPW or the g-formula. Using two concrete examples illustrated with causal diagrams, we demonstrate the limitations of existing graphical rules and their associated identification strategies for addressing selection bias.
@@ A.3 Simple graphical rules under SWIG by Mathur and Shpitser
Graphical rules based on SWIGs were also proposed for addressing selection bias [mathur2025simple]. It involves two sufficient conditions to analyze whether a conditional average treatment effect (CATE) can be identified by their formula.
Definition 3 (The graphical rules for assessing selection bias [mathur2025simple]).